### v3

# Why Vision Language Models Struggle with Visual Arithmetic? Towards Enhanced Chart and Geometry Understanding

Abstract
Vision Language Models (VLMs) have achieved remarkable progress in multimodal tasks, yet they often struggle with visual arithmetic, seemingly simple capabilities like object counting or length comparison, which are essential for relevant complex tasks like chart understanding and geometric reasoning. In this work, we first investigate the root causes of this deficiency through a suite of probing tasks focusing on basic visual arithmetic. Our analysis reveals that while pre-trained vision encoders typically capture sufficient information, the text decoder often fails to decode it correctly for arithmetic reasoning. To address this, we propose CogAlign, a novel post-training strategy inspired by Piaget’s theory of cognitive development. CogAlign trains VLMs to recognize invariant properties under visual transformations. We demonstrate that this approach significantly improves the performance of three diverse VLMs on our proposed probing tasks. Furthermore, CogAlign enhances performance by an average of 4.6% on Chocolate and 2.9% on Math-Vision, outperforming or matching supervised fine-tuning methods while requiring only 60% less training data. These results highlight the effectiveness and generalizability of CogAlign in improving fundamental visual arithmetic capabilities and their transfer to downstream tasks.  [^1] CogAlign data has been released at: https://github.com/SalesforceAIResearch/CogAlign.
Why Vision Language Models Struggle with Visual Arithmetic?
Towards Enhanced Chart and Geometry Understanding
Kung-Hsiang Huang1    Can Qin1    Haoyi Qiu2 Philippe Laban1    Shafiq Joty1    Caiming Xiong1    Chien-Sheng Wu1 1Salesforce AI Research    2UCLA 1{kh.huang, cqin, sjoty, cxiong, wu.jason}@salesforce.com
1 Introduction
In recent years, vision language models (VLMs) have rapidly advanced, demonstrating remarkable capabilities in integrating and processing multimodal information liu2023llava; dai2023instructblip; chen2024internvl; xue2024xgen. These models have found extensive applications across various domains, ranging from visual commonsense reasoning to sophisticated tasks like web agents xu2024llavacot; zhang2024improve; xie2024osworld; lin2024showui. By leveraging both visual and textual data, VLMs promise a nuanced understanding that surpasses what can be achieved by analyzing them individually.
Despite these advancements, current VLMs exhibit noticeable deficiencies in performing fundamental visual arithmetic: these models struggle with seemingly simple tasks like accurately counting objects, comparing lengths, assessing angles, and evaluating relative sizes or areas Rahmanzadehgervi_2024_blind; wang2024vdlm; huang2024frompixels; ullman2024illusion; wei2024slow; kamoi2024visonlyqa. These shortcomings are particularly evident in complex tasks such as chart understanding huang-etal-2024-lvlms and geometric problem-solving gao2023gllava.
In this study, we first delve into the root causes of VLMs’ difficulties with visual arithmetic, exploring several hypotheses to elucidate why VLMs often fail when faced with such challenges (Section 2). We propose a suite of probing tasks, focusing on basic visual arithmetic such as length comparison, to answer this question. Our analysis reveals that pre-trained vision encoders coupled with a simple linear classifier perform poorly on these probing tasks, indicating that a single linear layer is insufficient to decode the complex visual representations for arithmetic reasoning. However, when we fine-tune the text decoder of a VLM on these tasks, performance significantly improves. This suggests the bottleneck lies in the decoder’s ability to effectively process and utilize the visual information, rather than in the visual representation itself.
To tackle these challenges, we propose a novel post-training strategy, CogAlign, designed to improve the performance of VLMs in visual arithmetic tasks (Section 3). Drawing inspiration from Piaget’s theory of cognitive development piaget1952origins, our method focuses on enhancing VLMs’ understanding of conservation (recognizing that certain properties remain unchanged despite transformations) and decentration (considering multiple aspects simultaneously). We train VLMs using synthetically generated image pairs that demonstrate transformations, enabling them to compare and evaluate based on specific properties like length, angle, and quantity. By employing Direct Preference Optimization (DPO) rafailov2023dpo, the model learns from both positive and negative examples, offering a richer learning signal than traditional Supervised Fine-Tuning (SFT). Our experiments show that CogAlign significantly enhances performance across three VLMs of different scales and architectures on the proposed probing tasks.
Figure 1: Examples of probing tasks designed to assess visual arithmetic abilities. Each task presents a visual input and a question requiring comparison or evaluation of geometric properties. At the bottom of each task, we see that even top-performing VLMs like GPT-4o and InternVL2.5-78B struggle with these seemly simple tasks.
Furthermore, we evaluate CogAlign on two downstream benchmarks: Chocolate huang-etal-2024-lvlms for chart understanding, and Math-Vision wang2024mathv for geometric problem-solving (LABEL:sec:exps). Our results demonstrate the effectiveness of CogAlign in enhancing performance on these complex tasks. On average, CogAlign boosts performance by 4.6% and 2.9% on Chocolate and Math-Vision respectively, demonstrating that improving fundamental visual arithmetic capabilities translates to improved performance on downstream tasks. Notably, CogAlign outperforms or achieves comparable performance to SFT methods while requiring 60% less training data, even though CogAlign does not involve direct optimization for specific tasks. This showcases its strong generalizability and highlights its potential of focusing on foundational skills to unlock broader capabilities in VLMs.
Our main contributions are as follows:
  • We conduct an in-depth analysis to uncover the root causes of VLMs’ underperformance in tasks that involve visual arithmetic.
  • We develop CogAlign, a post-training strategy designed to enhance VLMs’ abilities in understanding performing visual arithmetic.
  • Extensive experiments on three VLMs show that CogAlign significantly improves performance in chart comprehension and geometric problem-solving, highlighting its generalizability.
2 Why Vision Language Models Struggle with Visual Arithmetic?
As suggested in previous studies, VLMs struggle with visual arithmetic Rahmanzadehgervi_2024_blind; wang2024vdlm, leading to poor performance in tasks involving such capabilities such as chart understanding huang-etal-2024-lvlms and geometric problem-solving gao2023gllava. In this section, we aim to understand the root causes behind such phenomenon. We first propose a suite of probing tasks we design to facilitate our analysis (Section 2.1) and then illustrate the various analyses we conduct to validate our hypotheses (Section 2.2).
Vision Encoder Angle Comparison Perpendicular Detection Length Comparison Chart Projection LLaVA-v1.5-proj 89.7 87.3 74.4 61.3 CLIP ViT-L/14 95.8 92.2 81.0 60.0 SigLIP-SO400M/14 98.5 92.1 89.9 74.5 InternViT-300M-V2.5 88.2 85.5 70.8 59.0 DINOv2-Large 96.2 94.7 81.8 57.3 Random guessing 50.0 50.0 50.0 50.0
Table 1: Accuracy (%) of different vision encoders with a linear classifier on the test set of each probing task. We conduct feature probing experiments by freezing the vision encoder and only fine-tuning the linear layer for binary classification. LLaVA-v1.5-proj refers to the representations obtained from the projection layer of LLaVA-v1.5.
2.1 Probing Tasks
We propose four probing tasks for assessing visual arithmetic capabilities, motivated by the fundamental operations needed to interpret visual data quantitatively. For a VLM to successfully understand a chart, for example, it must be able to compare lengths of bars or lines, discern relationships indicated by line slopes, and projecting points onto axes. An overview of the probing tasks are shown in Figure 1. All four tasks are discriminative and can be considered binary classification tasks. Below, we illustrate these tasks in details.
Angle Comparison
asks models to determine whether the angle of two wedges are the same. This requires the model to differentiate and measure angular magnitude, a seemly more complex operation that tests the model’s grasp of spatial relationships and angular geometry. This task assesses the model’s capacity to interpret rotational dimensions and engage in deeper analytical processing to distinguish subtle differences in angle, thereby evaluating the core geometric understanding of the model in angular perception.
Perpendicular Detection
challenges models to determine if two given lines are perpendicular to each other. Building upon the concept of angles, this task requires a deeper understanding of specific angular relationships, where perpendicularity implies a $90^{\circ}$ angle. While Angle Comparison focuses on general angle differentiation, Perpendicular Detection assesses a model’s ability to recognize this specific geometric configuration.
Length Comparison
asks models whether two lines with arbitrary slopes are of the same lengths. In addition to basic spatial reasoning, this task requires models to consider trigonometric relationships between the lines, demanding higher-level understanding of equivalence regardless of orientation. The variability in slopes necessitates an advanced ability to rotate or translate lines, challenging the model’s proficiency in geometric reasoning beyond simple horizontal and vertical comparisons.
Chart Projection
challenges the model to determine if the value of a red dot on a black line chart lies between 60 and 70. As the most complex task, this task integrates key aspects of the preceding tasks. It requires spatial reasoning to project the dot’s position onto the y-axis, similar to Angle and Perpendicular Detection. It then involves comparing the projected value’s magnitude against the specified range, akin to Length Comparison.
2.2 Probing Analysis
The research question we aim to answer is: Do visual representations from pre-trained vision encoders contain enough information to perform visual arithmetic tasks? To answer this question, we conduct experiments by feeding the outputs from various encoders into a linear classifier to perform binary classification on the probing tasks. For each task, we randomly generate 12,000 images programmatically with a train:development:test split of 10:1:1. Each split has a balanced portion of positive and negative labels. We test a wide range of vision encoder, including CLIP ViT-L/14 Radford2021LearningTV, SigLIP-SO400M/14 Zhai_2023_siglip, InternViT-300M-V2.5 chen2024internvl, and DINOv2-Large oquab2024dinov.We also evaluate the features produced by the projection layer of LLaVA-v1.5 liu2023llava. Each model (i.e., the single classifier) was trained for 200 epochs and the checkpoint that achieves the highest performance on the development set is selected.
The results are presented in Table 1. Overall, we observe that fixed visual representations, when paired with a single linear layer, yield reasonable performance on simpler tasks such as Simple Length Comparison and Angle Comparison. However, they struggle significantly with more complex tasks like Length Comparison and Chart Projection. Therefore, we conclude that pre-trained vision encoders do not convey sufficient information through their fixed visual representations for a linear classifier to succeed at visual arithmetic. This may be attributed to two potential reasons: (1) the visual representations genuinely lack the information necessary for visual arithmetic tasks, or (2) a linear layer lacks the capacity to effectively leverage the visual features provided.
To further investigate the underlying cause of this limitation, we perform additional experiments by fine-tuning the LLM-based text decoder component of LLaVA-v1.5, while keeping its vision encoder frozen. In VLMs, visual representations are concatenated with text representations in the decoder. Unlike a linear layer, the text decoder can process textual queries as inputs, offering an opportunity to understand the effect of textual clues provided by input queries. We test with three different queries: an Original query reflecting the task as shown in Figure 1, an Empty query which is a blank string, and an Irrelevant query such as “My name is John?”. Additionally, we evaluate LLaVA-v1.5 in a zero-shot setting for comparisons. Given our previous observations, these experiments focus exclusively on the Length Comparison and Chart Projection tasks.
VLM Query Type Fine-tuned? Length Comparison Chart Projection LLaVA-v1.5-7B - 50.0 51.3 Original 95.4 98.9 Empty 95.2 98.1 Irrelevant 95.8 97.8
Table 2: Accuracy (%) of fine-tuned LLaVA-v1.5-7B with different queries on the test set of each probing task. We conduct experiments by freezing the vision encoder and only fine-tune the LLM decoder for binary classification. The top row displays the zero-shot performance.
The fine-tuned LLaVA results are shown in Table 2. We have the following observations. First, existing VLMs do struggle with challenging visual arithmetic when used in zero-shot manners, achieving less than 90% and 75% on the two more challenging probing tasks, even with extensive in-domain training. The finding is consistent with prior studies Rahmanzadehgervi_2024_blind; wang2024vdlm and highlights the validity and complexity of our proposed probing tasks. Second, VLMs fine-tuned on in-domain data perform reasonably well in visual arithmetic. LLaVA-v1.5-7B is able to achieve an accuracy of above 95% on both Length Comparison and Chart Projection tasks. Third, the high performance of fine-tuned VLMs on in-domain data is due to the larger capacity of an LLM. Comparing the three different queries, we see the performance on Length Comparison and Chart Projection does not vary too much. This means that a fine-tuned LLaVA-v1.5-7B performs well even when the query provides no clue or irrelevant information about the given tasks. Combing this observation with our findings in Table 1, we learn that fine-tuned VLMs perform well because the LLM-based text decoder have larger capacity than a linear layer rather than leveraging the semantics of the input query.
2.3 Visual Representation Inspection
While the analysis conducted in Section 2.2 reveals important findings, one may argue that the fine-tuning success might be caused by the fine-tuned model exploiting patterns in binary tasks. Motivated by the visual representation analysis of rudman2025forgotten, we conduct additional experiments focusing on visual representations’ ability to differentiate isolated geometric properties. We generated datasets of images, each containing a single angle (sampled in 10-degree increments from 10° to 100°) or a single line (sampled in 0.1 length increments from 0.1 to 1.0). We created 100 images per class, resulting in 1,000 images each for the angle and line datasets. We then extracted visual embeddings for these images using the pre-trained CLIP and SigLIP vision encoders. To assess whether embeddings for different angles or lengths occupy distinct regions in the feature space (a prerequisite for successful differentiation or classification), we performed t-SNE visualization vandermaaten08tsne to understand cluster separation. The results are shown in Figure 2.
We have the following observations. First, there are some clear clusters for both CLIP and SigLIP embeddings, particularly for smaller angles and shorter lines. However, the separation is less distinct for longer lines and larger angles, especially with CLIP embeddings. Second, the separation for angles is more obvious than that for lines, aligning with our findings in Table 3 that Length Comparison is more challenging than Angle Comparison. Third, SigLIP embeddings demonstrate superior clustering for both angles and lines. This is also consistent with our Table 3 results where SigLIP outperforms CLIP across the proposed probing tasks.
Based on these observations and our findings in Section 2.2, we conclude that while visual representations from VLMs do encode information for visual arithmetic, it cannot be effectively decoded without further fine-tuning, which may in turn affect their zero-shot performance on downstream tasks like chart understanding and that fine-tuning success is not merely exploiting patterns in a binary task.
(a) Angle clustering by CLIP
(b) Line clustering by CLIP
(c) Angle clustering by SigLIP
(d) Line clustering by SigLIP
Figure 2: t-SNE visualizations of clustering by CLIP and SigLIP for angles and lines.
3 CogAlign
Figure 3: Example training data for CogAlign. Each example consists of a visual input, a query prompting comparison of a specific property (i.e. angle, length, distance, and etc), a positive response consistent with the visual input, and a negative response that contradicts it.
To address the challenges VLMs face in performing visual arithmetic, we propose a novel post-training method inspired by Piaget’s theory of cognitive development piaget1952origins, which outlines four stages: Sensorimotor, Preoperational, Concrete Operational, and Formal Operational. Each stage represents a different ability to process information and solve problems, culminating in abstract reasoning. The Concrete Operational Stage is particularly relevant. At this stage, children develop (1) conservation, understanding that certain properties like length remains constant despite changes in appearance, and (2) decentration, the ability to consider multiple aspects of a situation at once. These skills are essential for VLMs to perform visual arithmetic accurately, recognizing invariant properties such as length or angle across transformations. Current VLM training paradigms often neglect these cognitive processes, resulting in models that struggle to maintain key properties during visual transformations and to integrate multiple visual features effectively. While pre-trained visual encoders use losses that encourage some invariance to transformations, the integration of vision and language representations in decoders often lacks explicit enforcement of conservation and decentration principles, leading to models that capture visual features but fail to reason about them effectively.
To address these issues, we present a post-training method, Cognitive Alignment (CogAlign), aimed at enhancing VLMs’ understanding of conservation and decentration. Our approach explicitly trains VLMs to recognize invariant properties like length, angle, and count across different visual transformations. We achieve this by presenting the model with pairs of figures and associated queries designed to highlight these properties. The queries prompt the model to compare and contrast the figures, focusing on whether a specific property is different or same despite variations in appearance. This approach encourages the model to develop a stronger understanding of geometric concepts and move beyond superficial visual comparisons. Furthermore, we leverage DPO rafailov2023dpo for training, rather than SFT. DPO allows the model to learn from both positive and negative examples within the preference framework, providing a richer learning signal compared to SFT. By strengthening these cognitive capacities within VLMs, our goal is to improve their performance on tasks involving visual arithmetic. The subsequent subsections detail the specific training procedure employed (Section 3.1) and the automated construction of our training data (Section 3.2).
3.1 DPO Training Objective
Our goal is to train a model with parameters $\theta$ that learns conservation and decentration from contrasting responses by maximizing the conditional probability of positive responses over their negative counterparts. Concretely, the DPO training data consists of preference pairs, each containing a user query $Q$, an input image $I$, a positive response $R_{p}$ and a negative response $R_{n}$. The entire set of DPO training data can be represented as $\mathcal{D}=\left\{\left(Q,I,R_{p},R_{n}\right)^{(i)}\right\}_{i=1}^{\lvert\mathcal{D}\rvert}$. The objective function $\mathcal{L}_{\mathrm{DPO}}$ that DPO minimizes is:
$\mathcal{L}_{\mathrm{DPO}}(\pi_{\theta};\pi_{\mathrm{ref}})=-\mathrm{E}_{(Q,I,R_{p},R_{n})\sim\mathcal{D}}\Big{[}\log\sigma(r_{\Delta})\Big{]},$
$r_{\Delta}=\beta\log\frac{\pi_{\theta}(R_{p}|Q,I)}{\pi_{\mathrm{ref}}(R_{p}|Q,I)}-\beta\log\frac{\pi_{\theta}(R_{n}|Q,I)}{\pi_{\mathrm{ref}}(R_{n}|Q,I)},$
where $\sigma$ is the sigmoid function, $\pi_{\theta}$ is the parameterized policy under training, $\pi_{\mathrm{ref}}$ is the initial frozen policy, and $\beta$ is a hyper-parameter that controls the deviation from $\pi_{\mathrm{ref}}$.
3.2 Training Data Synthesis
To effectively train VLMs on the principles of conservation and decentration, we require a training dataset designed to highlight these concepts. This section details our automated process for synthesizing training data, encompassing visual generation and tailored query-response construction. We draw inspiration from our probing tasks but adapt the format to better suit the DPO training procedure. By plotting two shapes within a single image, we allow the model to directly compare invariant properties like length and angle across various transformations. We devise eight fundamental tasks, each of which aim to enhance VLMs’ different abilities to reason about visual arithmetic operations: understanding angle, length, distance, quantity, volume, position, slope, and intersection. An overview of these tasks is shown in Figure 3